# Open-Source Molecular Processing Pipeline for Generating Molecules

### v2

@@ 5 Conclusion & Discussion
In this work, we improve DeepChem’s generative modeling tools and provide a more standardized and scalable implementation that makes generative molecular methods more accessible to scientists. Standard Machine Learning practices are built within DeepChem (e.g., checkpointing, validation, logging, etc.), which would otherwise need some form of human expertise. Benchmarks show comparable performance with existing implementations, and the tight integration with DeepChem facilitates fast future improvements. This will also allow for Reinforcement Learning (Cao and Kipf, 2022) based approaches to be incorporated easily. We will provide multi-GPU training support in future work using Distributed Data Parallelism (Li et al., 2020), Sharding & PyTorch (Paszke et al., 2019).
@@
Appendix A MolGAN Pipeline
@@
1from deepchem.models.torch_models import BasicMolGANModel as MolGAN
2from deepchem.models.optimizers import ExponentialDecay
3import deepchem as dc
4import torch.nn.functional as F
5import torch
6...
7gan = MolGAN(learning_rate=ExponentialDecay(0.001, 0.9, 5000))
8dataset = dc.data.NumpyDataset([x.adjacency_matrix for x in features],[x.node_features for x in features])
9def iterbatches(epochs):
@@
19# train model
20gan.fit_gan(iterbatches(8), generator_steps=0.2, checkpoint_interval=0)
@@
1generated_data = gan.predict_gan_generator(10)
2# convert graphs to RDKitmolecules
3new_mols = feat.defeaturize(generated_data)
Appendix B NormalizngFlows Pipeline
@@
1import deepchem as dc
2from deepchem.data import NumpyDataset
3from torch.distributions.multivariate_normal import MultivariateNormal
4from rdkit import Chem
5from dc.torch_models.nflows import *
@@
7# Pass data through pipeline (mentioned above)
8...
@@
10# Construct flow model
11flows = [ActNorm(latent_size)]
12nfm = NormFlow(flows, distribution, dim)
13nfm.fit(max_iterations, optimizer)
@@
1mols = nfm.generate(num_molecules=1000)
2valid_mols = validate_mols(mols)

### v3

@@ 5 Conclusion & Discussion
In this work, we improve DeepChem’s generative modeling tools and provide a more standardized and scalable implementation that makes generative molecular methods more accessible to scientists. Standard Machine Learning practices are built within DeepChem (e.g., checkpointing, validation, logging, etc.), which would otherwise need some form of human expertise. Benchmarks show comparable performance with existing implementations, and the tight integration with DeepChem facilitates fast future improvements. This will also allow for Reinforcement Learning [2] based approaches to be incorporated easily. We observed high variance for datasets where the average number of atoms in a molecule was much larger than 9, and leave this for future work. We will provide multi-GPU training support in future work using Distributed Data Parallelism [14], Sharding & PyTorch [20].
@@
Appendix
Appendix A Mathematical Formulations
A.1 Molecules as Graphs
For MolGAN, each molecule can be represented as an undirected Graph $G$ with a set of edges $E$ and nodes $V$. Each atom $v_{i}\in V$ is associated with a $D$-dimensional one-hot vector $\mathbf{x}_{i}$. Each edge $(v_{i},v_{j})\in E$ is also associated with a bond type $y\in\{1,\dots,Y\}$. Thus, we have a representation of a graph as two objects: a $\mathbf{X}=[\mathbf{x}_{1},\dots,\mathbf{x}_{n}]^{T}\in\mathbb{R}^{N\times D}$ and an adjacency tensor $\mathbf{A}\in\mathbb{R}^{N\times N\times Y}$. $\mathbf{A}_{i,j}\in\mathbb{R}^{Y}$ is a one-hot vector indicating the type of edge between $i$ and $j$.
A.2 Generator
For any latent variable
$z\sim\mathcal{N}(\mathbf{0,I})$
the generator $\mathcal{G}_{\theta}(z)$ outputs two continuous objects:
Using the Gumbel Softmax trick [9], categorical sampling is performed as:
$\mathbf{\tilde{X}}$ $=\mathbf{X}+\text{Gumbel}(\mu=0,\beta=1)$ (2)
$\mathbf{\tilde{A}}$ $=\mathbf{A}_{ijy}+\text{Gumbel}(\mu=0,\beta=1)$ (3)
Alternatively, $\mathbf{\tilde{X}}$ and $\mathbf{\tilde{A}}$ can remain as $\mathbf{X}$ and $\mathbf{A}$.
A.3 Discriminator
The discriminator updates node representations through relational graph convolution as:
where $\boldsymbol{h}_{i}^{(l)}$ represents the signal at node $i$ and layer $l$, and $f_{s}^{(l)}$ is a self-connection transformation. After several layers of graph convolution, the graph score is calculated as:
where $\sigma$ is the sigmoid function $\sigma=\frac{1}{1+e^{-x}}$.
A.4 Normalizing Flow Model
The Normalizing Flow model starts with a base distribution $p_{z}$ and transforms it into a density function $p_{x}$. The transformation $x=L(z)$ is invertible, and the probability density is updated as:
$\rho_{X}(x)=\rho_{Z}(z)\cdot|det(J_{L}(z))|^{-1}$
where $J_{L}(z)$ is the Jacobian determinant of the transformation.
Appendix B Code Snippets
B.1 MolGAN Pipeline
@@
1 from deepchem.models.torch_models import BasicMolGANModel as MolGAN
2 from deepchem.models.optimizers import ExponentialDecay
3 import deepchem as dc
4 import torch.nn.functional as F
5 import torch
6 ...
7 gan = MolGAN(learning_rate=ExponentialDecay(0.001, 0.9, 5000))
8 dataset = dc.data.NumpyDataset([x.adjacency_matrix for x in features],[x.node_features for x in features])
9 def iterbatches(epochs):
@@
19 # train model
20 gan.fit_gan(iterbatches(8), generator_steps=0.2, checkpoint_interval=0)
@@
1 generated_data = gan.predict_gan_generator(10)
2 # convert graphs to RDKitmolecules
3 new_mols = feat.defeaturize(generated_data)
B.2 NormalizngFlows Pipeline
@@
1 import deepchem as dc
2 from deepchem.data import NumpyDataset
3 from torch.distributions.multivariate_normal import MultivariateNormal
4 from rdkit import Chem
5 from dc.torch_models.nflows import *
@@
7 # Pass data through pipeline (mentioned above)
8 ...
@@
10 # Construct flow model
11 flows = [ActNorm(latent_size)]
12 nfm = NormFlow(flows, distribution, dim)
13 nfm.fit(max_iterations, optimizer)
@@
1 mols = nfm.generate(num_molecules=1000)
2 valid_mols = validate_mols(mols)
Appendix C Hyperparameters
For our MolGAN implementation, similar to [2], we vary the given hyperparameters, and pick the best for each dataset prioritising Validity of generations. Except for QM9, we train each model for 30 epochs on the whole dataset. QM9 is trained similar to  [2], i.e., 300 epochs on a random 5k subset, we also perform early stopping if the uniqueness goes below 2%. Apart from this we fix Generator Steps to 0.2, i.e., one step every 5 discriminator steps, a batch size of 32, learning rate of $41e-$ 4, and an embedding dimension of 32.
Hyperparameter Values
Dropout Rates $\{0.0,0.1,0.25\}$
Sampling Mode Straight-through, Gumbel, SoftMax
Table 3: Hyperparameter settings for MolGAN.
For the Normalizing Flows implementation, we use a learning rate of 1e-4, and an equal weight decay. We train on a batch size of 1024 for 100 epochs. Both the models were trained on with Adam optimizer [10].